OpenStudy (anonymous):

for the given function, find the points on the graph at which the tangent line has slope 1. \[y=\frac{ 1 }{ 3 }x^3-2x^2+4x+1\]

OpenStudy (p0sitr0n):

derive the function and solve for the derivative = 1

OpenStudy (anonymous):

so would it just be y'=x^2

OpenStudy (p0sitr0n):

ye, so i guess its +1 and -1

OpenStudy (anonymous):

but that is not one of my choices. my choices are A. (1,3) and (3,4) B. (0,3) and (3,3) \[C.(1,\frac{ 10 }{ 3 }) and (3,4)\] \[D. (0,\frac{ 10 }{ 3 })and (3,4)\]

OpenStudy (anonymous):

\[y=-\frac{1}{3}x^3-2x^2+4x+1\\ y'=-x^2-4x+4\]

OpenStudy (anonymous):

ok so I understand where \[y=\frac{ 1 }{ 3 }x^3-2x^2+4x+1\] = \[y'=x^2-4x+4\] but how do I get the points from there?

OpenStudy (anonymous):

got it thanks the answer is C

OpenStudy (anonymous):

Yeah, you were supposed to let y' = 1, like @P0sitr0n said.